OpenStudy (heyitszia1):

PLEASE HELP I WILL FAN AND MEDAL! Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

OpenStudy (heyitszia1):

I only need help with part 3 :)

OpenStudy (heyitszia1):

I used \[2\sqrt{x+1}+4 = 2\] for extraneous and \[2\sqrt{x+1}+2 = 4\] for non extraneous.

OpenStudy (heyitszia1):

Can you help me @pooja195 @Luigi0210 @Cardinal_Carlo @satellite73 @Victoriasushchik

OpenStudy (heyitszia1):

@math&ing001

OpenStudy (heyitszia1):

Can you help me?

OpenStudy (math&ing001):

I'm guessing it's because a square root cannot be negative and in the first equation you have √(x+1)=-1

OpenStudy (heyitszia1):

@math&ing001 Thanks :)

OpenStudy (math&ing001):

Welcome =)